OpenStudy (studygurl14):

I DON'T KNOW HOW TO DO THIS! MEDAL! @jigglypuff314 @dan815 @jim_thompson5910

OpenStudy (anonymous):

http://gamutech.com

OpenStudy (studygurl14):

Please do not psam @joinusosisgoingdown

OpenStudy (studygurl14):

I think it is 3/5, but I'm not sure.... 50%+30%-20%....?

OpenStudy (dan815):

50 - work 30-children 20-both so 30 only work 10 only children 30+10=40 only work and children 40+20 = work, children and both so 60/100 = 3/5

OpenStudy (studygurl14):

thanks